A plane heads east with a veloncity of52m/s through a12m/s cross wind blowing the plane south find the magnitude and direction of the planes resultant veloncity (relative to due east)


1
Expert's answer
2021-02-17T10:56:04-0500


We can find the magnitude of the plane's resultant velocity from the Pythagorean theorem:


"R=\sqrt{v_p^2+v_w^2}=\sqrt{(52\ \dfrac{m}{s})^2+(12\ \dfrac{m}{s})^2}=53.4\ \dfrac{m}{s}."

We can find the direction of the plane's resultant velocity from the geometry:


"\theta=cos^{-1}(\dfrac{v_p}{R})=cos^{-1}(\dfrac{52\ \dfrac{m}{s}}{53.4\ \dfrac{m}{s}})=13^{\circ}\ S\ of\ E."